A 45.0 kg girl is standing on a 150 kg plank. The plank, originally at rest, is free to slide on a frozen lake, which is a flat, frictionless surface. The girl beginsto walk along the plank at a constant velocity of 1.50 m/s. What is her velocity relative to the surface of the ice?
Answer #1

mass of the girl, m1 = 45 kg
mass of the plank, m2 = 150 kg
initial velocity of the girl with respect toplank, U1 = 1.5 m/s
initial velocity of the plank , U2 = 0
final velocity of the ( plank + girl ) = V ( inopposite direction )
m1U1 + m2 U2 = ( m1 + m2 ) V
45* 1.5 + 0 = 150 V
V= 0.45m/s ( in opposite direction )
(a)
Velocity of girl relative to ice = U1 - V = 1.5- 0.45 = 1.05 m/s
(b)
Velocity of the plank, V = 0.45 m/s(opposite direction )

answered by: Jazmin